# Fast Subspace Fluid Simulation with Temporal-Aware Basis

### v1

@@ Abstract.
Through experimental validation on challenging scenarios, such as colliding vortex rings and boundary-interacting plumes, our method also exhibits superior performance and fidelity with significantly fewer basis functions compared to existing spatial ROMs. The inherent linearity of the DMD operator enables unique application modes, such as time-reversible fluid simulation. This work establishes another avenue for developing real-time, high-quality fluid simulations, enriching the space of fluid simulation techniques in interactive graphics and animation.
Simulation, Fluids, Smoke, Model-Reduction
ccs: Computing methodologies Modeling and simulationccs: Computing methodologies Physical simulation
Figure 1. Accurate Koopman-based Fluid Prediction of Vortex Ring Colliding. We compare (above) the ground truth MacCormack + Reflection (MC+R) solver (Zehnder et al., 2018) to (below) our reduced Koopman-based reconstruction, which leverages a reduced-order temporally-aware subspace with $r=150$ basis functions. Our approach achieves efficient, near-perfect prediction of arbitrary future fluid states via the $k$-th power of the eigenvalue matrix. In this experiment, parameter $k=100$.
[…]
@@ 1. Introduction
Figure 2. Long-time Single Step Integration. We demonstrate that our method can perform integration into arbitrary points in time via a single matrix multiply. Since the DMD operator is diagonal within the reduced basis, it is trivial to find the matrix that evolves the initial velocity field to the field at any point in time, significantly accelerating the integration as compared to traditional methods required by PCA.
[…]
@@ 2. Related Works
Following pioneering work by Foster and Metaxas (1996), Stam (1999a), and Fedkiw et al. (2001), the graphics community has made significant progress in the visual simulation of fluids. Our method straddles literature from both fluid control and runtime acceleration. From foundational work, we aim to demonstrate that DMD arises naturally as a combination of spatial reduced-order modelling and spectral decompositions, a unique position that makes it well suited for fluid control applications.
@@ 2.4. Spectral Methods
A key strength of spectral methods is the physical intuition of the basis functions. Whereas spatial order reductions construct basis functions based on the domain shape and are thus more physically intuitive for non-advecting phenomena such as elastic modes (Brandt and Hildebrandt, 2017; Sellán et al., 2023), spectral methods construct basis functions based on how the field evolves over time, leveraging the periodicity of fluid flows to represent the system’s dynamics in the frequency domain. Each mode thus corresponds to the propagation of a different wave, or coherent groups of waves. Kim et al. (2008) leveraged this to add smaller-scale turbulence as a post-processing step, and Chern et al. (2017) used this to directly modify vortex rings. This type of fluid control greatly motivates much of our application domains presented in Section 6.
@@ 2.5. Dynamic Mode Decomposition
DMD is a data-driven method that approximates the Koopman operator, a linear operator that describes the temporal evolution of a dynamical system (Schmid, 2010). Prior usage of DMD for fluids application has been limited, and have not been explored at all in computer graphics and creative applications. Usage has primarily been limited to playback and short-term forecasting (Proctor et al., 2016). Outside of fluid simulation, DMD has been used to process large datasets in high-dimensional systems (Williams et al., 2015), reducing noise in training data (Askham and Kutz, 2018), and constructing usable operators from noisy data (Sashidhar and Kutz, 2022). This latter extension, called OptDMD, is of significant interest to us, as the turbulence inherent in fluid fields naturally produces noisy training data.
@@
Figure 3. Reconstruction of 3D Plume Simulations. From left to right: standalone plumes, plumes interacting with a sphere, and plumes interacting with a bunny. For each configuration, the last frame of the original MacCormack (Selle et al., 2008) fluid simulation is shown on the left, alongside the last frame of our subspace simulation with SVD rank ranging from $r=2$ to $r=61$. Remarkably, the fluid dynamics demonstrate strong resilience to low-rank bases, highlighting a key advantage of our proposed reduced-order pipeline. Additionally, these scenarios illustrate the robustness of our method in handling increasingly complex boundary conditions. The reconstruction quality improves as the number of basis functions increases, enabling more accurate capture of finer details around the boundaries. While reasonable results are achieved with a small basis (r=9), increasing the basis significantly enhances the fidelity of the simulations. More details on the temporal evolution of the flow and other basis configurations can be found in the Fig. 10, Fig. 11 and Fig. 12.
[…]
3. Koopman Operator and Dynamic Mode Decomposition For Fluids
Consider an autonomous system of the form,
(1) $\frac{d}{dt}\bm{u}(t)=\bm{f}(\bm{u}(t))$
where $t$ is time, $\bm{u}\in\mathbb{R}^{N}$ is a state vector and $\bm{f}:\mathbb{R}^{N}\to\mathbb{R}^{N}$ is a nonlinear operator. There exists a time-dependent family of flow maps $\{\bm{F}_{t}\}_{t>0}$ such that $\bm{u}(t)=\bm{F}_{t}(\bm{u}_{0})$. That is, it maps some initial state $\bm{u}_{0}$ from time 00 to its state at time $t$. This induced flow map defines a time-series of data,
(2) $\bm{U}=\underbrace{\left[\begin{array}[]{c c c c c}|&|&&|\\ \bm{u}(0)&\bm{u}(1)&\cdots&\bm{u}(T)\\ |&|&&|\\ \end{array}\right]}_{T+1\;\text{Frames}}\begin{array}[]{l}\left.\begin{array}[]{c}\\ \\ \\ \end{array}\right\}\text{ $N$ DoFs}\end{array}$
which represent measured states of the system with their associated timestamps.
The finite-dimensional nonlinear trajectories induced by $\bm{f}(\bm{u})$ can be represented as infinite-dimensional linear trajectories of some scalar observables $g:\mathbb{R}^{N}\to\mathbb{C}$ in Hilbert space (Koopman, 1931). With these observables, Koopman (1931) defines a family of Koopman operators $\bm{\mathcal{K}}_{t}$ such that
(3) $\bm{\mathcal{K}}_{t}g=g\circ\bm{F}_{t}=g(\bm{u}(t)).$
That is to say, $\bm{\mathcal{K}}_{t}$ maps the measurement $g$ operator to its state at time $t$. Notice then, that we can recover our state $\bm{u}(t)$ by using the adjoint $g^{*}$ of our observable:
(4) $g^{*}\bm{\mathcal{K}}_{t}g=\bm{u}(t),$
Thus, if we can construct $\{\bm{\mathcal{K}}_{t}\}$ and $g$, we can get our system state at any point in time. We simply need to find the generator, given by:
(5) $\bm{\mathcal{K}}g=\frac{d}{dt}g\circ\bm{F}_{t}=\lim_{t\to 0}\frac{1}{t}(\bm{\mathcal{K}}_{t}g-g)$
Notice now that the generator $\bm{\mathcal{K}}$ is linear by construction, but acting on the space of nonlinear observables $g$. In essence, it relegates the nonlinearity to the extra dimensionality afforded by the function space traversed by $g$. This generator is typically what is called as The Koopman operator (as opposed to the family of operators $\{\bm{\mathcal{K}}_{t}\}$).
We additionally point out that if $\bm{\mathcal{K}}_{t}$ is a Koopman operator acting on an observable $g$, we can likewise refer to $\hat{\bm{\mathcal{K}}_{t}}=g^{*}\bm{\mathcal{K}}_{t}g$ as a Koopman operator acting on the observable $\bm{u}$. $g$ then is simply an operator that transforms observables of one function space (our state space) to another. For a finite dimensional vector $\bm{u}$, we represent the discretized Koopman operator as $\bm{K}$.
@@ 3.1. Approximating the Koopman Operator with Dynamic Mode Decomposition
We remind the reader that the Koopman operator acts on the infinite-dimensional function space spanned by some observables $g$. This, evidently, is cumbersome to work with. Thus, similar to the idea of applying PCA to reduce the dimensionality of a dataset in previous work (Kim and Delaney, 2013; Treuille et al., 2006), we instead approximate the Koopman operator using Dynamic Mode Decomposition (DMD) (Schmid, 2010), constructing a reduced, low-dimensional basis to represent the function space spanned by $g$.
Given two snapshots, a discrete Koopman operator is defined by:
@@
[…]
Intuitively, this optimizes ${{\bm{K}}}$ to minimize the difference between the next states $\bm{u}(t+\Delta{}t)$ and its prediction of the next states ${\bm{K}}\bm{u}(t)$.
@@
where $\bm{X}=\bm{U}\bm{\Sigma}\bm{V}^{T}$ is the singular value decomposition (SVD) of the snapshot matrix $\bm{X}$. We can truncate this eigensystem, taking the top $r$ singular values corresponding vectors to form the truncated Koopman operator $\tilde{\bm{K}}$. We can then project this truncated operator to the reduced basis $\bm{U}^{T}\tilde{\bm{K}}\bm{U}=\bm{U}^{T}\bm{X}^{\prime}\bm{V}\bm{\Sigma}^{-1}\bm{U}^{T}\bm{U}=\hat{\bm{K}}$ to produce the reduced Koopman operator $\hat{\bm{K}}$.
@@
Since applying the truncated Koopman operator $\tilde{\bm{K}}$ on the full space has complexity $\mathcal{O}(N^{2})$, we project the full-space velocity field $\bm{u}$ onto the reduced space spanned by the basis $\bm{\Phi}$:
@@
where $\bm{\Phi}^{+}=\bm{\Phi}^{*}(\bm{\Phi}\bm{\Phi}^{*})^{-1}\in\mathbb{R}^{r\times N}$ is the Moore-Penrose pseudoinverse of $\bm{\Phi}$ and $\bm{z}(t)=\bm{\Phi}^{+}\bm{u}(t)\in\mathbb{R}^{r}$ is the reduced state of the fluid system at time $t$. This projection step reduces the complexity to $\mathcal{O}(Nr)$.
@@
We note that $\bm{\Lambda}$ is the diagonal matrix of eigenvalues of $\hat{\bm{K}}$, and $\bm{\Phi}$ are the corresponding eigenvectors. That is to say, $\tilde{\bm{K}}=\bm{\Phi}\bm{\Lambda}\bm{\Phi}^{*}$ is exactly the spectral decomposition of the reduced Koopman operator.
@@
We would like to point out that both $\bm{\Lambda}$ and $\tilde{\bm{K}}$ are Koopman operators. $\bm{\Lambda}$ in particular gives particular insight into the theory; notice here that it acts on the function space of $g=\bm{\Phi}$. Each eigenfunction is thus a special observable, which clearly behaves linearly by definition. Since we know that $\bm{\Phi}^{*}$ maps observables to observables, $\tilde{\bm{K}}$ must necessarily also be a Koopman operator, this time acting on the identity operator (or more accurately, the $\bm{u}$ observable in Hilbert space).
The $\bf\Lambda$ representation in particular is particularly useful for us as it exposes the modes in an easily manipulatable manner. As a Koopman operator, it represents the time evolution of some observable. By being a diagonal matrix of eigenvalues, it turns out that these observables are exactly spatial modes that rotate with a particular frequency given by the imaginary part of the eigenvalues.
@@ 4.1. Constructing Training Data
We point out that the system solved by Equation 6 is agnostic to the structure of the input state $\bm{u}$. As stated, the DMD operator learns a mapping that evolves the state $\bm{u}$ from one point in time to its state at a future moment, and the natural description of a fluid state is to represent $\bm{u}$ as a stacked vector of fluid velocities, which implies that the DMD operator learns the PDE that governs the evolution of the fluid velocities over time.
@@
where $\bm{u}$ is the velocity field, $p$ is the pressure, $\rho$ is the density, and $\bm{g}$ is the gravitational force. We discretize the domain $\Omega$ with the staggered grid, a standard approach in computer graphics (Harlow et al., 1965), where the velocity field $\bm{u}$ is defined at the cell faces and the pressure $p$ is defined at the cell centers. This thus provides snapshots of the fluid velocity field $\bm{u}(t)$, at any point in time $t$. Note that $\bm{u}$ is a vector of size $\mathbb{R}^{N}$. This vector space henceforth will be our fullspace, as it includes spatially every point in our full resolution simulation.
Since our training data can be generated from any simulation method (or even observed from the real world), we utilize several fluid simulation algorithms to construct our dataset. Specifically, we include the standard stable fluids (Stam, 1999b), the MacCormack (Selle et al., 2008) (MC), the MacCormack + Reflection (Zehnder et al., 2018) (MC+R) and the Lattice Boltzmann Method with the Bhatnagar-Gross-Krook collision model (Chen and Doolen, 1998) (LBM-BGK).

### v2

# Fast Subspace Fluid Simulation with a Temporally-Aware Basis

@@ Abstract.
Through experimental validation on challenging scenarios, such as colliding vortex rings and boundary-interacting plumes, our method also exhibits superior performance and fidelity with significantly fewer basis functions compared to existing spatial ROMs. Leveraging the inherent linearity of the DMD formulation, we demonstrate a range of diverse applications. This work establishes another avenue for developing real-time, high-quality fluid simulations, enriching the space of fluid simulation techniques in interactive graphics and animation.
Model-Reduction, Koopman Theory, Dynamic Mode Decomposition
copyright: acmlicensedjournal: TOGjournalyear: 2025journalvolume: 44journalnumber: 4publicationmonth: 8doi: 10.1145/3730826ccs: Computing methodologies Modeling and simulationccs: Computing methodologies Physical simulation
Figure 1. Accurate Koopman-based Fluid Prediction of Vortex Ring Colliding. We compare (above) the ground truth MacCormack + Reflection (MC+R) solver (Zehnder et al., 2018) to (below) our reduced Koopman-based reconstruction, which leverages a reduced-order temporally-aware subspace with $r=150$ basis functions. Our approach achieves efficient, near-perfect prediction of arbitrary future fluid states via the $k$-th power of the eigenvalue matrix. In this experiment, parameter $k=100$.:
Figure 2. Long-time Single Step Integration. We demonstrate that our method can perform integration into arbitrary points in time via the exponential integration of a single matrix (Sec. 4.4). Since the DMD operator is diagonal within the reduced basis, it is trivial to find the matrix that evolves the initial velocity field to the field at any point in time, significantly accelerating the integration as compared to traditional methods required by PCA. Although the DMD operator is trained over the velocity field $\bm{u}$, we show the corresponding density field at each time point for clearer visualization. :
@@ 2. Related Works
Following pioneering work by Foster and Metaxas (1996), Stam (1999), and Fedkiw et al. (2001), the graphics community has made significant progress in the visual simulation of fluids. Our method straddles literature from both fluid control and runtime acceleration. From foundational work, we aim to demonstrate that DMD arises naturally as a combination of spatial reduced-order modelling and spectral decompositions, a unique position that makes it well suited for fluid control applications.
@@ 2.4. Spectral Methods
A key strength of spectral methods is the physical intuition of the basis functions. Whereas spatial order reductions construct basis functions based on the domain shape and are thus more physically intuitive for non-advecting phenomena such as elastic modes (Brandt and Hildebrandt, 2017; Sellán et al., 2023), spectral methods construct basis functions based on how the field evolves over time, leveraging the periodicity of fluid flows to represent the system’s dynamics in the frequency domain. Each mode thus corresponds to the propagation of a different wave, or coherent groups of waves. Kim et al. (2008) leveraged this to add smaller-scale turbulence as a post-processing step, and Chern et al. (2017) used this to directly modify vortex rings. This type of fluid control greatly motivates much of our application domains presented in Sec. 6.
@@ 2.5. Dynamic Mode Decomposition
DMD is a data-driven method that approximates the Koopman operator, a linear operator that describes the temporal evolution of a dynamical system (Schmid, 2010). Prior usage of DMD for fluids application has been limited, and have not been explored at all in computer graphics and creative applications. Usage has primarily been limited to playback and short-term forecasting (Proctor et al., 2016). Outside of fluid simulation, DMD has been used to process large datasets in high-dimensional systems (Williams et al., 2015) and construct usable operators from noisy data (Askham and Kutz, 2018). This latter extension, called OptDMD, is of significant interest to us, as the turbulence inherent in fluid fields naturally produces noisy training data.
@@
3. Koopman Operator and Dynamic Mode Decomposition
Consider a system of the form,
(1) $\frac{d\bm{u}}{dt}=\bm{f}(\bm{u}),$
where $\bm{u}(\bm{x},t)\in\mathbb{R}^{N}$ is a time-varying state vector in $\bm{x}\in\mathbb{R}^{d},\ t\in\mathbb{R}$, and $\bm{f}:\mathbb{R}^{N}\to\mathbb{R}^{N}$ is a nonlinear operator. Since the right-hand side is not explicitly dependent on time, this autonomous system evolves according to the current state $\bm{u}$, not when $\bm{u}$ is evaluated.
Suppose we collect a finite dictionary of observables, and stack them to yield the measurement vector,
(2) $\bm{g}(\bm{u}(t))=\begin{bmatrix}g_{1}(\bm{u}(t))\\ g_{2}(\bm{u}(t))\\ \vdots\\ g_{n}(\bm{u}(t))\end{bmatrix}=\begin{bmatrix}\bm{u}(\bm{x}_{1},t)\\ \bm{u}(\bm{x}_{2},t)\\ \vdots\\ \bm{u}(\bm{x}_{n},t)\end{bmatrix}\in\mathbb{R}^{Nn},$
where $n$ is the total spatial degrees of freedom of our measurements, and we take $T\leq n+1$ independent measurements of $\bm{g}(\bm{u})$ at different times. There exists an operator $\bm{K}$ on this basis that applies the following linearized time-stepping operation (Brunton et al., 2022),
(3) $\bm{g}_{k+1}=\bm{K}\bm{g}_{k},$
where the subscripts $\{k\in\mathbb{N}\ |\ 0\leq k<T\}$ indicate the time level of a given measurement. $\bm{K}$ is known as the restricted Koopman operator, and is guaranteed to exist and exactly recover the given input measurements as long as the number of time samples $T$ is fewer or equal to $n+1$, where $n$ is the size of the spatial degrees of freedom. Moving forward, for practical use, we only require the understanding that $\bm{K}$ exists, and we defer discussion of Koopman theory on why this is true to App. A.
@@ 3.1. Approximating the Koopman Operator with Dynamic Mode Decomposition
The restricted Koopman operator $\bm{K}$ has a basis with dimension on the number of degrees of freedom of the data (i.e. the number of grid points). In most reasonable simulation resolutions, this is computationally intractable, considering $\bm{K}$ is almost certainly dense in this grid basis.
To address this, we follow a strategy similar to prior dimensionality reduction approaches such as PCA (Kim and Delaney, 2013; Treuille et al., 2006), and instead approximate the Koopman operator using Dynamic Mode Decomposition (DMD) (Schmid, 2010), constructing a reduced, low-dimensional basis to represent the function space spanned by the input observables. Note that while $\bm{K}$ will exactly reproduce input measurements (and reproduce the continuous system in Eq. (1) in a least-squares sense on the finite measurement basis), the dimensionality reduction offered by DMD serves to approximate $\bm{K}$, and thus will introduce extra error in exchange for computational tractability.
As illustrated on the right, a key distinction between PCA and DMD is their relationship to time. Here, two consecutive snapshots are shown as two black dots (states recorded by frames) connected by a timeline. Conventional PCA is data order insensitive as its singular-value decomposition depends only on the covariance of the snapshot set, so re-ordering the frames leaves the low-rank surface where the projected states (red dots) reside unchanged.
DMD, in contrast, fits an operator $\bm{K}$ that advances the earlier state to the next predicted state (hollow red dot), as indicated by the red arrow. Because this least-squares fit relies on aligned pairs of samples, shuffling the data alters $\bm{K}$. This built-in temporal sensitivity lets DMD capture system dynamics rather than merely the geometry of the solution manifold, making it better suited for modelling time-evolving systems. Intuitively, PCA models a subspace of the solution space, whereas DMD models the operator that maps a state to the next—an objective that naturally requires temporal information.
Given two time-shifted sequence of snapshots $\bm{X}$ and $\bm{X^{\prime}}$, a discrete Koopman operator is defined by:
@@
$\text{where }\;\bm{X}$ $=\underbrace{\begin{bmatrix}\bm{u}(0)&\bm{u}(1)&\cdots&\bm{u}(T-1)\\ \end{bmatrix}}_{T\;\text{Frames}},$
$\bm{X}^{\prime}$ $=\underbrace{\begin{bmatrix}\bm{u}(1)&\bm{u}(2)&\cdots&\bm{u}(T)\\ \end{bmatrix}}_{T\;\text{Frames}},$
Intuitively, this optimizes $\bm{K}\in\mathbb{R}^{(Nn)\times(Nn)}$ to minimize the difference between the next states $\bm{u}(t+\Delta{}t)$ and its prediction of the next states ${\bm{K}}\bm{u}(t)$ under a linear model.
@@
where $\bm{X}=\bm{U}\bm{\Sigma}\bm{V}^{T}$ is the singular value decomposition (SVD) of the snapshot matrix $\bm{X}$, with $\bm{V}\in\mathbb{R}^{(Nn)\times r}$ being the left eigenvectors, $\bm{\Sigma}$ being the eigenvalues matrix, and $\bm{U}\in\mathbb{R}^{(Nn)\times r}$ being the right eigenvectors. We can truncate this eigensystem, taking the top $r$ singular values corresponding vectors to form the truncated Koopman operator $\tilde{\bm{K}}$. We can then project this truncated operator to the reduced basis $\bm{U}^{T}\tilde{\bm{K}}\bm{U}=\bm{U}^{T}\bm{X}^{\prime}\bm{V}\bm{\Sigma}^{-1}\bm{U}^{T}\bm{U}=\hat{\bm{K}}$ to produce the reduced Koopman operator $\hat{\bm{K}}$.
@@
Since applying the truncated Koopman operator $\tilde{\bm{K}}$ on the full space has complexity $\mathcal{O}(N^{2}n^{2})$, we project the full-space velocity field $\bm{u}$ onto the reduced space spanned by the basis $\bm{\Phi}$:
@@
where $\bm{\Phi}^{+}=(\bm{\Phi}^{*}\bm{\Phi})^{-1}\bm{\Phi}^{*}\in\mathbb{R}^{(Nn)\times r}$ is the Moore-Penrose pseudoinverse of $\bm{\Phi}$ and $\bm{z}(t)=\bm{\Phi}^{+}\bm{u}(t)\in\mathbb{R}^{r}$ is the reduced state of the fluid system at time $t$. This projection step reduces the complexity to $\mathcal{O}(Nnr)$.
@@
We note that $\bm{\Lambda}$ is the diagonal matrix of eigenvalues of $\hat{\bm{K}}$, and $\bm{\Phi}$ are the corresponding eigenvectors. That is to say, $\tilde{\bm{K}}=\bm{\Phi}\bm{\Lambda}\bm{\Phi}^{*}$ is exactly the spectral decomposition of the reduced Koopman operator. This low-rank model allows linear simulation of the reduced state, offering an efficient surrogate for the nonlinear underlying system.
@@
We would like to point out that despite both $\bm{\Lambda}$ and $\tilde{\bm{K}}$ are are referred to as Koopman operators in the literature, we remark that in our formulation they act on a subspace of linear observables derived from state components. Therefore, the dynamics we capture are strictly linear, and our model is best interpreted as a linear approximation of the full underlying nonlinear system.
$\bm{\Lambda}$ in particular gives particular insight into the theory; notice here that it acts on the function space of $g=\bm{\Phi}$. Each eigenfunction is thus a special observable, which clearly behaves linearly by definition. Since we know that $\bm{\Phi}^{*}$ maps observables to observables, $\tilde{\bm{K}}$ must necessarily also be a Koopman operator, this time acting on the identity operator (or more accurately, the $\bm{u}$ observable in Hilbert space).
In addition, the $\bf\Lambda$ representation exposes the modes in an easily manipulatable manner. As a Koopman operator, it represents the time evolution of some observable. By being a diagonal matrix of eigenvalues, it turns out that these observables are exactly spatial modes that rotate with a particular frequency given by the imaginary part of the eigenvalues.
Figure 3. Reconstruction of 3D Plume Simulations. From left to right: standalone plumes, plumes interacting with a sphere, and plumes interacting with a bunny. For each configuration, the last frame of the original MacCormack (Selle et al., 2008) fluid simulation is shown on the left, alongside the last frame of our subspace simulation with SVD rank ranging from $r=2$ to $r=61$. Remarkably, the fluid dynamics demonstrate strong resilience to low-rank bases, highlighting a key advantage of our proposed reduced-order pipeline. Additionally, these scenarios illustrate the robustness of our method in handling increasingly complex boundary conditions. The reconstruction quality improves as the number of basis functions increases, enabling more accurate capture of finer details around the boundaries. While reasonable results are achieved with a small basis (r=9), increasing the basis significantly enhances the fidelity of the simulations. More details on the temporal evolution of the flow and other basis configurations can be found in the Fig. 10, Fig. 11 and Fig. 12.:
@@ 4.1. Constructing Training Data
We point out that the system solved by Eq. (4) is agnostic to the structure of the input state $\bm{u}$. As stated, the DMD operator learns a mapping that evolves the state $\bm{u}$ from one point in time to its state at a future moment, and the natural description of a fluid state is to represent $\bm{u}$ as a stacked vector of fluid velocities, which implies that the DMD operator learns the PDE that governs the evolution of the fluid velocities over time.
@@
where $\bm{u}$ is the velocity field, $p$ is the pressure, $\rho$ is the density, and $\bm{g}$ is the gravitational force. We discretize the domain $\Omega$ with the staggered grid, a standard approach in computer graphics (Harlow et al., 1965), where the velocity field $\bm{u}$ is defined at the cell faces and the pressure $p$ is defined at the cell centers. This thus provides snapshots of the fluid velocity field $\bm{u}(t)$, at any point in time $t$. Note that $\bm{u}$ is a vector of size $\mathbb{R}^{Nn}$. This vector space henceforth will be our fullspace, as it includes spatially every point in our full resolution simulation.
Since our training data can be generated from any simulation method (or even observed from the real world), we utilize several fluid simulation algorithms to construct our dataset. Specifically, we include the standard stable fluids (Stam, 1999), the MacCormack (Selle et al., 2008) (MC), the MacCormack + Reflection (Zehnder et al., 2018) (MC+R) and the Lattice Boltzmann Method with the Bhatnagar-Gross-Krook collision model (Chen and Doolen, 1998) (LBM-BGK).